Circular No. 6544

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970111 A. Guarnieri, C. Bartolini, and A. Piccioni, University of Bologna;

N. Masetti, University of Padua; M. Teodorani, Osservatorio Astronomico, Napoli; E. Costa, Istituto Astrofisica Spaziale, Frascati; and D. Dal Fiume and F. Frontera, TESRE, Bologna, communicate: “A preliminary analysis of CCD observations of the GRB 970111 field (IAUC 6533), obtained with the Bologna University 1.5-m telescope in B and R bands (limiting magnitudes B ∼ 20, R ∼ 21) on Jan. 14.19, 15.19, and 17.19 UT, shows the presence of some objects that are probably variable; however, on this timescale, no object reveals a decline > 0.3 mag. A comparison between these frames and the same field from the Digital Sky Survey (DSS) shows that all the objects brighter than B ∼ 19 and R ∼ 20 in our frames are also present in the DSS field and that their luminosities have not noticeably changed since then (1950.294). Many galaxies are present in the GRB 970111 field. We also observed the error-box fields of sources ‘a’ and ‘b’ (IAUC 6539) in the R band, and we detected five and eight objects, respectively, brighter than R ∼ 21. None of these seems to be variable in this time interval (or if compared with the DSS). Our group continues the optical observations of the field.”

CYGNUS X-1 R. P. Fender and C. Brocksopp, University of Sussex; and G. G. Poo-

ley, University of Cambridge, report: “We have detected the 5.6-day orbital modulation of the blackhole candidate Cyg X-1 in radio monitoring at 15 GHz with the Ryle Telescope (Cambridge), undertaken since 1996 Oct., during which time the source has been in a hard x-ray state. This confirms the apparent orbital modulation in previous Very Large Array data (Han and Hjellming, Ap.J., submitted). The modulation is approximately sinusoidal, with a semiamplitude of ∼ 3 mJy superimposed on a mean flux density of ∼ 14 mJy for the entire data set. Orbital flux minimum occurs at superior conjunction of the compact object, as observed also in ASM/XTE x-ray data (IAUC 6510). This implies that the radio-emitting region is associated with the compact object, and the orbital modulation may be due to varying line-of-sight optical depth through the stellar wind of the OB-type companion star. The orbital modulation is superimposed upon other trends that also appear to be correlated with the x-ray activity.”

1997 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green